A car starts moving rectilinearly first with acceleration f = 5 m/s2 (the initial velocity is zero) then uniformly moving for some time and finally deceleration at the same rate f and comes to rest. The total time of motion is 25 sec. the average velocity during the time equals to v = 20 m/s2. How long the car moves uniformly?

Grade:11

2 Answers

Vikas TU
6 years ago
Dear Student,
From the given data:
acceleration f = 5
      u=0
      t=t1
        
Distance=S1

B: a=0
     u1=5t1
       
t=t2
Distance=S2

C: a=-f=-5
      u2=5t1
      v=0
Distance=S1 (Since same deceleration and same change in velocity)
time=t1 (same time as A , as magnitude of decellaration is same as acceleration)

2t1+t2=25
Avg velocity=2S1+S2/25 = 20
=>2S1+S2=500m

for case A
S1=ut+at2/2
=>S1=5xt12/2
=>S1=v2/10

for case B
S2=vt2
=>S2=
 x t2

2S1+S2=500
=>v2/5 + vt2 = 500
=> v2/5 + v(25-2t1)=500
=>v2/5 +25v-2v2/5=500
=>v2-125v+2500=0
=>v=100m/s(cant be the answer because t2 would be negative) or 25m/s

t1=v/5=5s (20s cant be the answer because t2 would be negative)
t2=25-10=15s.
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)
Jatin
5 years ago
One must give clear cut answer to questions
.
It should not confuse people .
People like the clear cut ans and also like it